OpenStudy (anonymous):

A particle with an initial velocity of 12 m/s has its acceleration modeled by the equation a(t) = t 2 + 3. Find the particle’s velocity at t = 2

OpenStudy (anonymous):

a(2) = (2)^2 + 3 solve

OpenStudy (anonymous):

possible answers can only be 4 8.667 14.667 20.667

OpenStudy (anonymous):

Yes. My fault. Sorry. Too early for me. v(t) = integral of acceleration v(t) = v(0) + (1/3) t^3 + 3t evaluated at t=2 v(2) = 12 + (8/3) + 6 = 20.667

OpenStudy (anonymous):

thank you sooooooo much !! have a nice day !!

OpenStudy (anonymous):

You are very welcome. Have a great day, too.